OpenStudy (anonymous):

The length of a rectangle is five times its width. If the perimeter is at most 96 centimeters, what is the greatest possible value for the width? A. 40 cm B. 19.2 cm C. 16 cm D. 8 cm

OpenStudy (anonymous):

sorry i have to go for a bit

OpenStudy (lynfran):

perimeter=2(length*width) (p/2)=l*w (p/2)-l=w and length length=8 so (96/2)-8=w 48-8=w ....=w

OpenStudy (lynfran):

8*5=w...

OpenStudy (anonymous):

and i am back

OpenStudy (anonymous):

A